OpenStudy (anonymous):

the perimeter of a rectangle is 38 cm. If the length is one less than three times the width, what is the width?

OpenStudy (anonymous):

perimeter=2length+2width p=2l+2w and we know that l=3w-1 so we can replace l with 3w-1: p=2(3w-1)+2w p=6w-2+2w 38= 8w+2 36=8w 36/8=w 4.5=w

OpenStudy (anonymous):

thank you very muchh!! i was confused about one of the steps and now i'm clear with it!